OpenStudy (minatsukisaya95):

how do we integrate (y^(1/2))/(y+1)

OpenStudy (dmb1215):

what is this geometry

OpenStudy (anonymous):

Change the variable to \[ u=\sqrt y\\ du= \frac {dy}{2 \sqrt y}\\ dy =2 \sqrt y du= 2 u du\\ \]

OpenStudy (anonymous):

\[ \int \frac{\sqrt{y}}{y+1} \, dy=\int\frac{2 u^2 du}{u^2+1}=\\ \int\left(2-\frac{2}{u^2+1}\right)du= 2 u - 2 \tan^{-1}u + C\\ 2 \sqrt{y}-2 \tan ^{-1}\left(\sqrt{y}\right)+C \]

OpenStudy (minatsukisaya95):

thank u that helps me ;-)

OpenStudy (anonymous):

YW